Measured viral load in plasma of two patients are given in the following table Patient -1 (RNA copies/ml) Patient- 2 (RNA copies/ml) Days post infection 12 14 16 21 23 1353 10081 33612 706640 1097500 79232 798550 1680700
Assuming that the virus grows exponentially, use the given data to calculate initial viral load, growth rate, and doubling time for each patient. Show your best fit model and data in the graph for each patient. Compare viral dynamics in these two patients. a. b. C.
Answer #1

Hi.

Using Excel you can find the models by linearizing de exponential models

y=ax(1 + r)r

Where y is the viral load at a time x, a is the initial vital load and r is the growth rate

This can be calculated using a model like:

y=a\: b^x

and linarizing is you get:

Log(y) = Log(a) + x Log(b)

to find Log a and Log b you use least squares regression.

m = Log(b)

c=Log(a)=\frac{\sum Log(y)-m\sum x}{N}

a. We make the following table to find m and c for Patient 1

x days P1 Log(y) x*Log(y) x^2
14 10081 4.003504 56.04905 196
16 33612 4.526494 72.42391 256
21 706640 5.849198 122.8332 441
23 1097500 6.040405 138.9293 529
74 20.4196 390.2354 1422

m= Log(b)= 0.235 → b= 100. 235-1.718

c = Log(a) = 0.751 → a = 100.75-5.636

For Patient 1 the model is:

y 5.636 × 1.718

Meaning that the initial viral load is 5.636

The growth rate is 0.718

And the doubling time can be calculated by:

Log(2) Log(1.718) 1281days

We make the following table to find m and c for Patient 2

x days P2 Log(y) x*Log(y) x^2
12 1353 3.131298 37.57557 144
16 79232 4.898901 78.38241 256
21 798550 5.902302 123.9483 441
23 1680700 6.22549 143.1863 529
72 20.15799 383.0926 1370

m = Log(b)= 0.274 → b= 100. 24-1.879

cLog(a) 0.114a 10114-1.3

For Patient 2 the model is:

y=1.3\times1.879^x

Meaning that the initial viral load is 1.3 approx 6

The growth rate is 0.879

And the doubling time can be calculated by:

Log(2) Log(1.87可:J·099days = 1 ,099days 1.8791-2 → x=

b. Using desmos graphing calculator to graph the models we get

y-5.636x1.718 y = 1.3 x 1.879 150000 沓 mpulsado po desmos

c. From the graph you can see that even so the Patient 1 started with a higher initial viral load the growth rate is lower that for Patient 2, and with the over time patient 2 will eventually have a higher viral load because his initial viral load was lower but the growth rate was higher.

Note: i personally recomend that you get more meassurements, if you can twice a day, if you can get more data, you can have a more precise model to predict the behavior of the virus.

If you liked my answer please dont forget to give it a thumbs up, it would help me a lot. Best regards